The mean annual cost of autobile insurance is R95. Assume that the population standard deviation is R14. What is the probability that a simple random sample of size 30 for automobile insurance policies will have a sample mean less than R90?


A.2.5560 B. 0.0250 C. -1.96

D.0.0205

E.0.9750


1
Expert's answer
2021-07-20T06:46:15-0400

Let "X=" a sample mean: "X\sim N(\mu, \sigma^2/n)."

Given "\mu=95, \sigma=14, n=30"


"P(X<90)=P(Z<\dfrac{90-95}{14/\sqrt{30}})"

"\approx P(Z<-1.956152)\approx0.025224"

B. 0.0250